# Trust, Because You Can’t Verify: Privacy and Security Hurdles in Education Technology Acquisition Practices

### v1

@@ 7.1 Sharing liabilities during data breaches
Privacy incidents, such as data breaches or vendors misusing data, are becoming commonplace in the EdTech space [46, 3]. P9 identified the large number of EdTech tools per institution—ranging from hundreds to upwards of thousands of services—as one of the major causes of privacy issues:
@@
In many cases, HEIs make public announcements about the incident, because it is either legally required to give “specific notifications” (P3) based on the data lost or to control reputational damage. P2 stated:
@@
Contracts play a vital role in establishing vendors’ obligations regarding public notification, legal recourse, damage assessment, and the prevention of future incidents in the event of privacy violations. As P5 noted, “that’s [contract] the most serious leverage we can have.” Similarly, P1 told us that “if they [vendor] committed to maintaining data privacy, then that’s the contractual obligation, and there’s not much we can do from that standpoint.”
Collaborative Information Sharing. In the context of addressing privacy incidents, P3 described how they use their collaboration with other universities to help inform them of vendors issues and how to make vendors follow necessary obligations:
{quoting}
“We had an incident in the recent past [with data leakage] that we, let’s just say, leveraged the REN-ISAC that the community is sharing such that others were made aware of some of the issues of this third party. So not only do we leverage the contract and enforce upon the vendor their obligations to notify other institutions that may have been impacted and notify students that may have been impacted, but also leverage that community network that we have to help inform others so that they know the extent of what’s happened”
Legal Implications and Ambiguous Liability. Privacy incidents can carry significant legal repercussions and hefty fines, particularly if they involve protected data (e.g., health or financial data), as discussed by P1: {quoting}“With HIPAA data you also have to notify the federal government that there was a HIPAA data loss and you could be fined thousands of dollars per record for that loss.”
Determining liability for privacy-violating incidents can be challenging, particularly when considering whether the university or the affected vendor should be held accountable. This challenge is exacerbated if sub-vendors are involved, given that their contract with the primary vendor and the data flow between them is invisible to the university. P11 expressed their concern regarding this ambiguity:
@@
Takeaway: HEI vulnerability to vendor breaches Vendor breaches pose significant liabilities for HEIs, with contracts often leaving ambiguities and inadequate recourse once data is compromised. Moreover, ambiguities in federal regulations in data breach liability (e.g., as it relates to FERPA) limits HEI ability to hold vendors accountable.
@@ 7.2 Visibility and Transparency concerns
Lack of Auditability. The ability to perform or require third-party audits is a large part of the acquisition process, it allows for HEIs to develop a level of trust and transparency with technologies; However, once the EdTechs are acquired, institutions lose their ability to audit those technologies until the next licensing cycle.
@@
“We continue to refine the processes, but that is one of the product managers responsibility… [Asking] where’s that data going? Is the place that it is going, did they get a security evaluation? When was the last time they got a security evaluation? If somebody changes a vendor, until our contract gets renewed and we redo another evaluation, we’d never know that unless the vendor tells us. There’s no contractual trigger”
@@ 7.3 Refining HEIs’ processes and practices
Despite challenges, HEIs implement measures to keep their internal system secure and refine internal practices regularly.
Internal Audits and Penetration Tests. To prevent external tools from introducing vulnerabilities into HEIs’ internal systems, some institutes conduct annual internal audits and penetration tests (P5 and P6), especially with vendors that directly plug into their systems. P3 discussed it:
@@
HEIs rigorously review their practices as a crucial protective measure. P2 provided insights into their institution’s audit procedures, highlighting both the frequency and scope of their internal audits. They conduct periodic audits, with top systems reviewed every two to three years and ancillary systems every three to five years. P2 elaborated on the details of their internal audit process:
@@
“We have an internal policy that does at least two audits and then penetration tests yearly on internal systems and services, that does extend [to] the cloud. The audit will go through the Active Directory, which is our identity authorization directory and it might ask us how we do data security and backup. They’ll come through and ask industry standard questions: are we following the right processes, … the right controls.”
However, P2 expressed frustrations as such procedures are ‘unsustainable for checking every vendor that their institution has.” Meanwhile, P1 has taken proactive steps to mitigate this issue by strategically reducing the number of technologies employed and opting for solutions that integrate multiple capabilities. This helps delay the potential scalability problem, which is also highlighted by P3.
Adapting to Evolving Technologies. As part of continuously refining audit processes, some HEIs are considering artificial intelligence (AI) based tools but they also acknowledge associated risks. AI could help HEIs protect their own data and increase surveillance over vendor tools. P11 shared their thoughts on AI-enabled tools:
@@
Takeaway: Refining processes and practices While not having the ability to confirm that vendors are performing their obligations, HEIs focus on their own practices and looking ahead at the growing landscape of emerging tools such as AI that could assist or harm them.
@@ 7.4 Off-boarding EdTech
Takeaway: Off-boarding process When discontinuing an EdTech, HEIs have no power or ability to investigate if their data has been fully erased off of a vendor’s system. HEIs are made to trust a written verification made by the vendor itself, and even that is only sometimes provided.
8 Challenges in Managing HEI Data Post-acquisition (RQ4):
As HEIs navigate the post-acquisition complexities of EdTech integration, they are exposed to issues surrounding data ownership, off-boarding procedures, and recourse in the event of privacy breaches. We discuss our participants’ thoughts on these challenges here.
8.1 Off-boarding EdTech
EdTech discontinuation is common, prompting an important question: What happens to the data collected by the vendor? Common reasons for discontinuation include price increases (P1, P3, P6, and P8), vendor disputes (P5, P6, and P8), and the identification of better tools (P1, P2, and P8).
Regardless of the reason, when services are discontinued, vendors typically initiate an off-boarding process outlined in the contract, which ideally involves erasing all institutional data from their systems and providing a confirmation document. Vendors however, may not be as cooperative as one might like; P2 complained that:
{quoting}
“once you tell them [cloud vendors] you’re not buying anything from any of them anymore, they are much less inclined to help you off board… it might be as little as half of the off boardings we get the confirmation that. Yes, everything’s shut down and all [the university’s] data is gone”
While vendors may be contractually obliged to provide a “written assurance” (P5) about data erasure, universities need to pursue vendors for assurance, especially for sensitive data. P2 went on explaining:
{quoting}
“Unless it’s in the contract, unless we want to pursue long squabbles with a disinterested party, we essentially have to come to a conversation. If we had 200,000 of our alumni with all their Social Security numbers on it, you can bet we’re going to pursue it. If we have 200,000 people with their addresses and a phone number, we’re likely to pursue it. If we have 1000 people and it’s just their first name, we’re probably going to write it off.”
P12 emphasized that while a signed assurance provides “some legal protection,” it may not fully address concerns if data remains with the vendor despite claims of erasure. P5 shared that due to vendors’ non-cooperation, the off-boarding for complex systems could take multiple years. P3 highlighted a trend towards centralized systems managing various functions from “student information systems to HR to finance.” They expressed concerns if institutions wish to move away from such integrated systems:
{quoting}
Similar to data ownership, contractual language becomes key during off-boarding process, yet, concerns regarding vendor integrity emerge.
8.2 Handling Privacy Violations
Privacy incidents, such as data breaches or vendors misusing data, are becoming commonplace in the EdTech space [46]. P9 identified the large number of EdTech tools per institution—ranging from hundreds to an upwards of thousand services—as one of the major causes of privacy issues:
{quoting}
P2 shared that the proliferation of sub-vendors, with inadequate resources, exacerbates the risks of sub-vendor breaches.
{quoting}
In many cases, HEIs make public announcements about the incident, because it is either legally required to give “specific notifications” (P3) based on the data lost or to control reputational damage. P2 stated:
{quoting}
In the event of privacy violations, contracts play a vital role in establishing vendors’ obligations regarding public notification, legal recourse, damage assessment, and ensuring prevention of future incidents. However, as P5 noted “that’s [contract] the most serious leverage we can have.” Similarly, P1 told us that “if they [vendor] committed to maintaining the data privacy, then that’s the contractual obligation that there’s not much we can do from that standpoint.”
Collaborative Information Sharing. In the context of addressing privacy incidents, P3 described how they use their collaboration with other universities to help inform them of vendors issues and how to make vendors follow necessary obligations:
{quoting}
“We had an incident in the recent past [with data leakage] that we, let’s just say leveraged the REN-ISAC that the community is sharing such that others were made aware of some of the issues of this third party. So not only do we leverage the contract and enforce upon the vendor their obligations to notify other institutions that may have been impacted and notify students that may have been impacted but also leverage that community network that we have to help inform others so that they know the extent of what’s happened”
Legal Implications and Ambiguous Liability. Privacy incidents can carry significant legal repercussions and hefty fines, particularly if they involve protected data, as discussed by P11: {quoting}“the fines around healthcare records are really scary. Our CISO told us that if they captured all of our healthcare records, he thought it would be about a \$50 million fine.”
Determining liability in such incidents can be challenging, particularly when considering whether the university or the affected vendor should be held accountable. This challenge is exacerbated if sub-vendors are involved, given that their contract with the primary vendor and data flow between them is invisible to the university. P11 expressed their frustrations regarding this ambiguity:
{quoting}
“Is it the university [who is liable]? Is it the vendor of the university? Is it the sub vendor or the sub sub vendor? Who is liable? To date, the Department of Education and some very, very sketchy advice says ultimately the university runs the application, they’re liable.”
While contracts help to some extent in cases of privacy violations, the need for clearer guidelines and accountability mechanisms become crucial in navigating these situations.
@@ 9 Discussions and conclusions
Overall, our study suggests that HEIs privacy posture is largely underlined by the protections offered by FERPA. This result is consistent with Brown and Klein [22], who found that out of $151$ institutional policy documents, $81$ were FERPA-related. Given the limited data types that are protected by FERPA, the coverage of entities (only federally funded organizations), and the existence of loopholes that can be leveraged to evade compliance altogether, the privacy of institutional data is in jeopardy.
@@
Security assessments dilemmas. Even when compliance is mandatory, currently there is no way to guarantee it. HEIs mostly rely on a questionnaire-based assessment of EdTech before procurement (e.g., HECVAT [24]), without any actual visibility into the tools and auditing capabilities. Some HEIs even lack resources to examine those assessment reports in the face of a growing number of vendors. We note that assessments like HECVAT and SOC2 focus on securing the data from unauthorized access, but they do not cover how this data can be used by ‘authorized’ entities, and yet many vendors are unwilling to complete this basic assessment step.
@@
Contractual Challenges and Power Asymmetry. Contracts dictate data ownership and use, and we uncovered the challenges HEIs face to create contracts with concrete and privacy-focused terms and conditions regarding data collection and use. First, most HEIs lack dedicated personnel who can go back and forth multiple times with the vendors to review and revise the contract, and this situation is exacerbated by the continuous addition of new tools that are coming from different vendors. Second, even when HEIs try to address ambiguity in security assessments and contract documents or revise terms and conditions to reflect proper data use and accountability, the vendors push back. There is a huge power asymmetry between vendors and HEIs, even those who are among the largest public universities in the US. While some universities form coalitions, which helps to some extent in amending vendors, there are challenges in collaborations, and most universities and community colleges are left out. Third, there is an inherent challenge HEIs face in making concrete statements about data collection: HEIs cannot examine the tools themselves, and so it is incredibly difficult to verify from outside what these tools are doing, what data they are collecting, and where they are sending this data. Thus, HEIs are left to trust vendors, which is tricky as pointed out by some of our interviewees.
Who owns the data? Our study also uncovered grim realities regarding data ownership and use. While HEIs fight to keep ownership of the data, it may not be exclusive and may depend on how hard they can fight. Vendors can also become the sole owner upon service discontinuation since HEIs cannot ensure the deletion of their data. We speculate on other ways to own data: the contract allows vendors to use the data (even if they do not own it), and vendors can have sub-vendors with the same privilege. Sub-vendors can change at any time and have an arbitrary contract with the primary vendor that allows them to own/share the data with other parties (including the primary vendor) or data brokers without any restriction. Future studies could shed light on this sub-vendor dynamics.
@@
Obscurity of (improper) data utilization. Data can also create new data through inference [15]. For example, since behaviors correlate with demographics and other personal factors, they can be “inferred” from interaction data collected by many EdTech tools [15], allowing vendors to learn more about the students. Besides the issues of ownership of and consent for this new data, their use has profound ethical implications [47], and may disproportionately harm vulnerable groups (e.g., outing a non-binary student by “inferring” their gender [15]). Indeed, one of the primary uses of student data is creating such machine learning-based predictive models [15]. Companies like Instructure (that built Canvas), offer services based on these models [48], and their takeover by a for-profit equity farm further heightens the ethical implications of digitizing education [49].
@@
Risks and liabilities of AI/ML in EdTech landscape. We also learned that while (sub-)vendors profit from data by building and selling new (AI/ML-based) services, they don’t share the risks and liabilities. Vendors actively try to add terms that would waive their liability, supporting prior research [21], and even when they are in the contracts, the long chain of (sub-)vendors, where they can update their privacy policy anytime and without the knowledge of the HEIs, makes it nearly impossible to properly assign the liability during a privacy violation. Moreover, while HEIs may demand some level of protection for the data vendors collect, there is no provision for the protection of ML models trained from that data, and recent research shows that they are susceptible to attacks that can reveal the data they were trained with by querying or inverting the model [50].
Recommendations. We conclude with some recommendations. First, we strongly advocate for a comprehensive privacy law at the federal level so that HEIs would have some baseline protection regardless of how strongly they can negotiate with the vendors. We also recommend federal or state mandate standard contracts with by-default data protection clauses, including not only who can use data but also specifying the purposes. We invite researchers from different disciplines, including education, law, and computing, to collaborate in creating these standards. As these laws and standards will apply to all vendors, they will prevent privacy issues caused by long vendor chains or the deployment of EdTech without following the institutional procurement process. We end with optimism noting that some HEIs are forming data governance boards, we hope that security and privacy researchers will collaborate to identify appropriate privacy-enhancing tools, as well as research and develop new tools that meet the specific needs of education processes.

### v2

@@ 7.1 Sharing Liabilities during Data Breaches
Privacy incidents, such as data breaches or vendors misusing data, are becoming commonplace in the EdTech space [49, 3]. P9 identified a large number of EdTech tools per institution—ranging from hundreds to upwards of thousands of services—as one of the major causes of privacy issues:
@@
Contracts play a vital role in establishing vendors’ obligations, as P5 noted, “that’s [contract is] the most serious leverage we can have.” Similarly, P1 told us that “if they [vendor] committed to maintaining data privacy, then that’s the contractual obligation, and there’s not much we can do from that standpoint.” In many cases, HEIs make public announcements about the incident, because it is either legally required to give “specific notifications” (P3) based on the data lost or to control reputational damage. P2 stated:
@@
Legal implications and ambiguous liability. Privacy incidents can carry significant legal repercussions and hefty fines, particularly if they involve sensitive and protected data (e.g., health or financial data), as discussed by P1: {quoting}“With HIPAA data you also have to notify the federal government that there was a HIPAA data loss and you could be fined thousands of dollars per record for that loss.”
Determining liability for privacy-violating incidents can be challenging, particularly when considering whether the university or the affected vendor should be held accountable. This challenge is exacerbated if sub-vendors are involved, given that their contract with the primary vendor and the data flow between them is invisible to the university (§ 6.4). P11 expressed their concern regarding this ambiguity:
@@
Takeaway: HEI vulnerability to vendor breaches Vendor breaches pose significant liabilities for HEIs, with contracts often leaving ambiguities and inadequate recourse once data is compromised. Moreover, ambiguities in federal regulations in data breach liability (e.g., as it relates to FERPA) limit HEI’s ability to hold vendors accountable.
@@ 7.2 Visibility and Transparency Concerns
Lack of auditability. The ability to perform or require third-party audits is a large part of the acquisition process, it allows for HEIs to develop a level of trust and transparency with technologies (§ 6.3). However, once the EdTech are acquired, institutions lose their ability to audit those technologies until the next licensing cycle.
@@
“We continue to refine the processes, but that is one of the product managers responsibility… [Asking] where’s that data going? Is the place that it is going, did they get a security evaluation? When was the last time they got a security evaluation? If somebody changes a vendor, until our contract gets renewed and we redo another evaluation, we’d never know that unless the vendor tells us. There’s no contractual trigger.”
@@ 7.3 Refining HEIs’ Processes and Practices
Because HEIs cannot audit EdTech in operation, some of them have implemented measures to keep their internal system secure and regularly refine those practices.
Internal audits and penetration tests. To prevent external tools from introducing vulnerabilities into HEIs’ internal systems, some institutes conduct annual internal audits and penetration tests, especially with vendors that directly plug into their systems. P3 discussed it:
@@
HEIs rigorously review their practices as a crucial protective measure. P2 highlighted their institution’s periodic audit procedures, noting that top systems are reviewed every two to three years and ancillary systems every three to five years. P2 elaborated on the details of their internal audit process:
@@
“We have an internal policy that does at least two audits on internal systems and services, that does extend [to] the cloud. The audit asks us how we do data security and backup, ask industry standard questions: are we following the right processes, the right controls?”
However, P2 expressed frustrations that such procedures are “unsustainable for checking every vendor that their institution has.” To mitigate this, P1 and P3 take proactive steps to strategically reduce the number of technologies employed and opt for solutions that integrate multiple capabilities.
Adapting to evolving technologies. As part of continuously refining audit processes, some HEIs are considering AI tools, while acknowledging associated risks. AI could help HEIs protect their own data and increase surveillance over vendor tools. P11 shared their thoughts on AI-based tools:
@@
Takeaway: Refining processes and practices While not having the ability to confirm that vendors are performing their obligations, HEIs focus on their own practices and look ahead at the growing landscape of emerging tools such as AI that could assist or harm them.
@@ 7.4 Off-boarding EdTech
Takeaway: Off-boarding process When discontinuing an EdTech, HEIs have no power or ability to investigate if their data has been fully erased from a vendor’s system. HEIs are made to trust a written verification made by the vendor itself, and even that is only sometimes provided.
@@ 8 Discussions and conclusions
This study aims to surface specific security and privacy challenges faced by HEIs that come from the specific tools they use, the resources and processes they have in place, and HEI-specific policies. Thus, while some findings may apply to many HEIs, we do not aim for generalizability. Below, we discuss our results by highlighting key challenges and providing recommendations based on our study that can help strengthen security and privacy within HEIs regarding EdTech acquisition, retention, and discontinuation.
@@
Limitations of FERPA in safeguarding HEI. Overall, our study suggests that HEIs’ privacy posture is largely underlined by the protections offered by FERPA (§ 5.1). This result is consistent with Brown and Klein [20], who found that out of $151$ institutional policy documents, $81$ were FERPA-related. Given the limited data types that are protected by FERPA [17], the coverage of entities (only federally funded organizations) [17], and the existence of loopholes that can be leveraged to evade compliance altogether [50], the privacy of institutional data is in jeopardy.
@@
Security assessments dilemmas. Although compliance with FERPA and HEIs’ security practices is mandated [17], currently there is no way to guarantee it from the vendors’ end (§ 6.3). HEIs mostly rely on a questionnaire-based assessment of EdTech before procurement (e.g., HECVAT [22]), which does not provide any actual visibility into EdTech and its auditing capabilities. Some HEIs even lack resources to examine those assessment reports in the face of a growing number of vendors (§ 6.4). We note that assessments like HECVAT and SOC2 focus on securing the data from unauthorized access, but they do not cover how this data can be used by “authorized” entities. Alarmingly, many vendors are unwilling to complete even these basic assessment procedures, highlighting a critical gap in data protection that has significant implications for student privacy and institutional accountability.
@@
Contractual challenges and power asymmetry. Contracts dictate data ownership and use. However, we uncovered that HEIs face challenges to create contracts with concrete and privacy-focused terms and conditions regarding data collection and use (§ 6.2). First, the web of sub-vendors (§ 6.4) exacerbates the issues of ensuring robust data protection practices at the contract level. Second, even when HEIs try to address ambiguity in security assessments and contract documents or revise terms to reflect proper data use and accountability, vendors often push back. This reveals a significant power asymmetry between vendors and HEIs, including those among the largest public universities in the US. While some universities form coalitions to improve their negotiating position, these efforts face collaboration challenges and often exclude smaller universities and community colleges. Third, HEIs face inherent difficulties in making concrete statements about data collection since they cannot independently examine the tools, making it incredibly difficult to verify from outside what these tools are doing, what data they are collecting, and where they are sending this data. Consequently, this power imbalance and lack of transparency forces HEIs to trust vendors’ assurances, a practice fraught with risk, as highlighted by our participants.
Data ownership challenges. Having addressed the broader contractual challenges, our study also uncovered grim realities regarding data ownership and use in the context of EdTech acquisitions (§ 6.2). Firstly, HEIs strive to retain ownership of their data, but achieving exclusive control often depends on the institution’s negotiating power. Secondly, contracts may permit vendors to use HEI data without owning it, raising concerns about control and usage rights. Thirdly, vendors can grant similar data usage privileges to sub-vendors, who can change at any time and may have arbitrary contracts that allow them to share or sell data with other parties (including the primary vendor) or data brokers without any restriction. Lastly, vendors might gain sole ownership of data if services are discontinued, as HEIs lack mechanisms to ensure data deletion. The complex relationships of HEIs with vendors and sub-vendors underscore the necessity for further research to unravel the implications of these dynamics on data control and privacy in EdTech acquisitions.
@@
Obscurity of (improper) data utilization. Data also creates new data. For example, behaviors often correlate with demographics and other personal factors, which can be “inferred” from interaction data collected by various EdTech tools [13], allowing vendors to learn more about the students. In addition to deciding the ownership for this newly created data, their use has profound ethical implications [51] and the potential to disproportionately harm vulnerable groups (e.g., outing a non-binary student by “inferring” their gender [13]).
Indeed, one of the primary uses of student data is creating such machine learning-based predictive models [13]. Companies like Instructure, which developed Canvas, offer services based on these models [52]. When such companies are acquired by for-profit equity firms, as seen with Canvas [53], it further intensifies the ethical issues surrounding the digitization of education. The ethical standards of the previous company may not necessarily be upheld under new ownership, and the prospect of misuse or sale of extensive student data raises serious concerns [50, 4]. Educational technologists are increasingly raising alarms about such company takeovers and other issues with the EdTech ecosystem [10, 54].
@@
Risks and liabilities of AI/ML in EdTech landscape. We also learned that while (sub-)vendors profit from data by building and selling new (AI/ML-based) services, they don’t share the associated risks and liabilities (§ 7.3). Moreover, vendors often include terms that waive their liability, as also noted in prior research [19]. Even when liability terms are included in contracts, the complex chain of (sub-)vendors—who can update their privacy policies without HEIs’ knowledge—makes it nearly impossible to assign liability in the event of a privacy violation. Although HEIs may seek protection for the data collected by vendors, there are no safeguards for the ML models trained from that data. Notably, recent research shows that ML models are susceptible to attacks that can reveal the very data used to train them through model inversion or querying [55]. These emerging concerns necessitate a reevaluation and enhancement of regulatory frameworks.
Recommendations:
For regulators. We strongly advocate for comprehensive federal privacy laws to provide HEIs with baseline protection regardless of their negotiating power with vendors. We also recommend federal or state mandates for standard contracts with default data protection clauses that will apply to all (sub-) vendors to prevent privacy issues caused by complex vendor chains and non-compliant EdTech deployments.
For researchers. Significant research efforts are required to enable transparency, auditability, and accountability in the EdTech ecosystem. Researchers could develop non-intrusive tools to record and analyze data flows for oversight and auditing purposes, especially post-acquisition. Privacy researchers could study techniques to establish private-by-design paradigms, such as differential privacy for aggregated data, minimal purpose-specific data collection, and verifiable data deletion protocols. Legal scholars could scrutinize existing laws to identify policy gaps that vendors exploit and create more comprehensive frameworks benefiting HEIs and student privacy. Researchers in business and management could investigate the impact of company acquisitions on data policies and handling practices, particularly for large vendors like Canvas. Finally, policy researchers could develop methods to ensure vendors comply with data protection standards before and after procurement, enabling HEIs to be confident that vendors are not misusing or mishandling their data throughout the procurement process.
For HEIs. We recommend HEIs increase inter-HEI collaborations to exchange knowledge about best practices and experiences with vendors, which will help strengthen their data security and privacy posture. Forming coalitions will enable HEIs to be more confident in their vendor agreements and enforce better data protection practices. For example, by being upfront and asking for clarifications on vague clauses, supply chains (including third-party vendors, sub-vendors, shadow IT, etc.), and the off-boarding process. Additionally, involving cybersecurity and policy can help HEIs design rigorous vendor assessment processes, negotiate clear data use agreements, and conduct regular audits and compliance checks. HEIs could also adopt data minimization practices, advocate for stronger regulations, and leverage more third-party security assessments.
We end with optimism noting that some HEIs are forming data governance boards, we hope that security and privacy researchers will collaborate to identify appropriate privacy-enhancing tools, as well as research and develop new tools that meet the specific needs of education processes.